### v2

# AIvaluateXR: An Evaluation Framework for on-Device AI in XR with Benchmarking Results

Abstract
The deployment of large language models (LLMs) on extended reality (XR) devices has great potential to advance the field of human-AI interaction. In case of direct, on-device model inference, selecting the appropriate model and device for specific tasks remains challenging. In this paper, we present AIvaluateXR, a comprehensive evaluation framework for benchmarking LLMs running on XR devices. To demonstrate the framework, we deploy 17 selected LLMs across four XR platforms—Magic Leap 2, Meta Quest 3, Vivo X100s Pro, and Apple Vision Pro—and conduct an extensive evaluation. Our experimental setup measures four key metrics: performance consistency, processing speed, memory usage, and battery consumption. For each of the 68 model-device pairs, we assess performance under varying string lengths, batch sizes, and thread counts, analyzing the tradeoffs for real-time XR applications. We finally propose a unified evaluation method based on the 3D Pareto Optimality theory to select the optimal device-model pairs from the quality and speed objectives. Additionally, we compare the efficiency of on-device LLMs with client-server and cloud-based setups, and evaluate their accuracy on two interactive tasks. We believe our findings offer valuable insight to guide future optimization efforts for LLM deployment on XR devices. Our evaluation method can be followed as standard groundwork for further research and development in this emerging field. The source code and supplementary materials are available at nanovis.org/AIvaluateXR.html.
Index Terms:
Extended reality, large language models, human AI interactions, performance evaluation, conversational user interfaces
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
Extended Reality (XR) and Artificial Intelligence (AI) have become increasingly prominent, and their intersection is even more compelling for researchers and real-world applications [1, 2, 3, 4, 5, 6]. Since the 2022 release of OpenAI’s ChatGPT interface [7], relying on the GPT-3.5 Large Language Model (LLM), we have witnessed the disruptive and transformative effects of LLMs. These models are capable of describing a wide variety of topics, respond at various levels of abstraction, and communicate effectively in multiple languages. They have proven capable of providing users with accurate and contextually appropriate responses. In addition to general applications such as grammar and text processing [8, 9], automated chatbot services, and source code generation [10], LLMs are increasingly being utilized in computer vision, human-computer interaction, and visual understanding.
In computer vision, language models are combined with visual signals to perform tasks such as verbal scene description and open-world scene graph generation [11]. In user interaction and visualization research, LLMs serve as verbal interfaces for controlling software functionality or adjusting visualization parameters [12, 13]. Another disruptive application of LLMs involves their use in wearable technology [14] to assist users in various environments. For instance, Meta AI is being integrated into Meta’s virtual reality hardware, where images captured by the device are streamed over the network to Meta’s servers for model inference, enabling image interpretation or prompt response tasks [^1] https://www.meta.com/blog/quest/meta-ai-on-meta-quest-3/. The overarching vision is to develop assistive hardware that is as lightweight as sunglasses but capable of aiding users in a wide range of tasks by understanding the scene they are viewing.
While such technological advances have the potential to empower users in unprecedented ways, the need for on-device AI in XR arises in several scenarios: (i) regions with limited or no internet connectivity (e.g., field training or remote maintenance); (ii) strict security and privacy requirements (e.g., healthcare, defense); (iii) reducing network latency; (iv) enabling autonomous operation as both LLMs and XR hardware continue to advance rapidly; (v) avoiding subscription costs; and (vi) providing the ability to fine-tune models for specific tasks and domains. In this regard, several smaller LLMs have been released by Meta, Mistral AI, Microsoft, and Google [15]. These models are designed to operate efficiently on mobile devices and extended reality hardware, typically consisting of billions of trainable parameters, while some LLMs with fewer than a billion parameters have already demonstrated promising performance [15].
Local execution of LLMs on XR devices is expected to become increasingly necessary for a wide range of applications. However, selecting the optimal device and model for a specific application is a complex decision. Device specifications and model documentation alone are not sufficient for making this decision. Furthermore, the hardware configurations of XR devices (such as CPU, memory capacity, and thermal management) and the architectures of the LLMs can significantly influence the results. Therefore, defining standardized evaluation criteria that address these complexities and ensures fairness across tests is also a research challenge. To the best of our knowledge, no comprehensive study currently provides clear guidelines for such evaluations.
To address these challenges, we present AIvaluateXR, a comprehensive framework for deploying LLMs on XR platforms and systematically evaluating their performance across multiple metrics. AIvaluateXR works in two steps: first, it deploys LLMs on XR devices and evaluates their performance across a range of factors; then, in the second step, it applies Pareto analysis to the results to identify the best-performing model-device pairs. We applied AIvaluateXR to deploy 17 LLMs across four XR devices—Magic Leap 2, Meta Quest 3, Vivo X100s Pro [^2] Vivo X100s Pro is, strictly speaking, not an XR device. We include it in our comparison because smartphones are commonly used as XR platforms, and this device features competitive hardware specifications., and Apple Vision Pro—and conducted a thorough evaluation in terms of performance consistency (i.e., stability over time), processing speed, memory usage, and battery consumption. Finally, we used Pareto analysis to identify the best-performing model-device pairs among the 68 combinations evaluated.
Considering the multiple challenges and factors affecting the evaluation, we first identify key evaluation metrics, and then design an experimental setup to minimize potential bias. A key challenge is the effect of varying string lengths, batch sizes, thread counts, and background processes. To account for string length variation, we conducted experiments using different lengths. To mitigate hidden factors like background processes, each experiment was repeated five times, and average values were calculated. To provide task-specific analysis, experiments were conducted across all LLMs tasks: Prompt Processing (PP) for evaluating performance with varying prompt lengths, Token Generation (TG) for varying generated token set sizes, plus Batch Test (BT) and Thread Test (TT) for assessing parallelism and concurrency efficiency. We believe our experimental protocol and results will serve as a foundation for further research and development in this emerging field. In summary, this paper contributes:
  • An evaluation framework and an experimental setup for assessing the performance of LLMs on XR devices, which can be used as standard guidelines for future research.
  • A comprehensive evaluation of 17 LLMs on four XR devices, resulting in 68 model-device pairs, across four critical metrics: processing speed, memory usage, stability over time, and battery consumption.
  • The deployment of various LLMs on XR devices, along with an analysis of their accuracy for interactive applications using standard query datasets, and a comparison with cloud-based and client-server-based LLMs in terms of efficiency and accuracy.
2 Related Work
We review related work along two major research directions: LLM-powered XR and on-device LLMs. For further details on AI in XR, see the recent review articles [1, 2, 4, 3].
2.1 LLM-powered XR Applications
Owing to the powerful semantic understanding and extensive general knowledge of LLMs, numerous studies have explored their application in assisting various tasks within XR scenarios. Recent studies have demonstrated the use of LLMs for natural interaction and conversational interfaces in VR and AR scenarios [16, 17, 18]. LLMs have also been applied for content generation, chat-based assistance, VR education[19], social VR [20], and as intelligent companions in XR applications [21, 22]. In addition to LLMs, vision-language models (VLMs) have been integrated into AR systems to enhance visual and multimodal understanding [14], scene-driven AR storytelling [23, 24], contextual guidance, dynamic content creation, and much more [25].
XaiR [14] integrates Multimodal Large Language Models (MLLMs) with XR devices using a client-server architecture, where computationally intensive MLLM tasks are offloaded to a server while spatial context is handled locally on the XR headset. The system supports real-time multimodal input; however, latency—comprising both network delays and LLM inference time—remains a challenge, often requiring users to wait 4–6 seconds for a response. Torre et al.[26] introduced a similar framework, LLMR, which leverages LLMs for real-time creation and modification of interactive mixed reality experiences, enabling tasks such as generating new content or editing existing elements on VR/AR devices.
Jia et al. [12] developed the VOICE framework, which employs a two-layer agent system for conversational interaction and explanation in scientific communication, with a prototype deployable on VR devices. Kurai et al. [27] proposed MagicItem, a tool that allows users with limited programming experience to define object behaviors within the metaverse platform. Zhang et al. [28] introduced OdorAgent, which combines an LLM with a text-image model to automate video-odor matching. Yin et al. [29] identified potential limitations in LLM-based automated systems and proposed the systematic framework Text2VRScene to address them. Chen et al. [30] leveraged the extensive capabilities of LLMs in context perception and text prediction to enhance text entry efficiency by reducing manual keystrokes in VR scenarios. Giunchi et al. [31] developed DreamCodeVR, a tool designed to help users, regardless of coding experience, create basic object behavior in VR environments by translating spoken language into code within an active application. Wan et al. [32] presented an LLM-based AI agent for human-agent interaction in VR, involving GPT-4 to simulate realistic NPC behavior, including context-aware responses, facial expressions, and body gestures.
XR applications often require 3D content that can be generated using AI. For example, Kawka et al. [33] proposed a method to create 3D objects from textual and image data, along with techniques for surface remeshing and simplification. Other approaches include character animation [34], text-to-3D [35], speech-to-3D generations [36], and image-to-3D reconstruction [37]. Recent work has also explored modeling human avatars and motion [38], as well as spatial understanding and object recommendation [39]. Although these systems typically rely on cloud-based models or client-server architectures rather than performing inference directly on the device, they demonstrate a clear need for AI in XR. This dependence raises concerns about user privacy, latency, costs, and internet access requirements. In our work, we evaluate the local inference performance of various LLMs and propose a prototype capable of running LLM inference entirely on-device.
2.2 On-device LLMs
Running LLMs on edge devices, commonly called on-device LLMs, has garnered significant research interest due to their advantages in enhancing privacy, reducing latency, and operating without the need for internet connectivity. Because of the limited memory and computing capabilities, on-device LLMs usually require resource-efficient LLM deployment [40]: a trade-off between performance and model size.
Cheng et al. [41] introduced the SignRound method, which leverages signed gradient descent to optimize both rounding values and weight clipping. This approach achieves outstanding performance in 2- to 4-bit quantization while maintaining low tuning costs and eliminating additional inference overhead. Ma et al. [42] developed the LLM-Pruner method, which uses structural pruning to selectively remove non-essential coupled structures based on gradient information, effectively preserving the core functionality of the LLM. Their results demonstrate that the compressed models continue to perform well in tasks such as zero-shot classification and generation. Gu et al. [43] introduced a novel knowledge distillation method that compresses LLMs into smaller models, resulting in the student model MINILLM. MINILLM demonstrates superior performance compared to baseline models, producing more precise responses with reduced exposure bias, improved calibration, and enhanced long-text generation capabilities. Liu et al. [44] developed a tuning-free 2-bit KV cache quantization algorithm, named KIVI, which independently quantizes the key cache per channel and the value cache per token. This approach enabled up to a 4× increase in batch size, resulting in a 2.35×2.35$\times$2.35 × to 3.47×3.47$\times$3.47 × improvement in throughput on real LLM inference workloads. Liu et al. [45] introduced MobileLLM, which explores the importance of model architecture for sub-billion-parameter LLMs by leveraging deep and narrow architectures, combined with embedding sharing and grouped-query attention mechanisms.
Although these studies primarily focused on reducing model size efficiently for deployment on various devices, they did not evaluate the models’ performance in real-world on-device scenarios, which is the focus of our paper.
3 Research Methods and Materials
This section outlines our research methodology, including the key challenges, the LLM deployment approach, and the evaluation metrics and models used in this study.
Problem Statement: Given a set of LLMs =$\mathcal{M}=\{m_{i}\}_{i=1}^{N}$ and a set of XR devices $\mathcal{D}=\{d_{j}\}_{j=1}^{K}$, AIvaluateXR evaluates the performance of each model-device pair $(m_{i},d_{j})$ across several key factors and quantitative metrics. It first generates evaluation results for each metric, which are then passed to a Pareto analysis module to determine the Pareto-optimal set. We used AIvaluateXR to deploy 17 LLMs, denoted as $\{m_{1},m_{2},\dots,m_{17}\}$ (see Table I), across four XR and mobile devices: $d_{1},d_{2},d_{3},$ and $d_{4}$, representing Magic Leap 2 (ML2), [^3] https://www.magicleap.com/magicleap-2 Meta Quest 3 (MQ3), [^4] https://www.meta.com/quest Vivo X100 Pro (Vivo), [^5] https://www.vivo.com/en and Apple Vision Pro (AVP), [^6] https://www.apple.com/vision-pro respectively. Our goal is to perform a comprehensive evaluation of these models on the four devices to determine how well they handle large-scale language processing tasks in resource-constrained XR environments.
We define our evaluation using six key performance metrics: $p_{0},p_{1},p_{2},p_{3},p_{4},p_{5}$, which represent Model Quality, Performance Consistency, Processing Speed (w.r.t. string length and prompt length), Parallelism (w.r.t. thread count and varying batch size), Memory Usage, and Battery Usage, respectively. Fig. 1 presents an overview of our research pipeline, where each rectangular box represents a specific procedure, and the connected ovals indicate their respective outcomes. These are discussed in detail below.
3.1 LLM Deployment on XR Devices
We aim to deploy the LLMs locally on XR devices. For this purpose, we customize the Llama.cpp library [46] and build it for the four target XR devices. The resulting application is capable of loading various appropriately sized GPT-Generated Unified Format (GGUF) models. Our deployment leverages the functionalities of Llama, allowing us to run basic LLMs tasks, and adjust various parameters. For example, to evaluate different aspects in an experiment, we can control various parameters such as prompt length in Prompt Processing (PP) and the size of the token set in Token Generation (TG). To benchmark the models on various devices, the top-level scripts in Llama.cpp are compiled into binary executable files. These binaries include scripts specifically designed for testing both execution speed and model quality. For Apple Vision Pro, we utilized Xcode [^7] https://developer.apple.com/xcode/, Apple’s official IDE, to build and deploy testing scripts directly to the device. For the remaining three devices, the testing was performed via a shell interface opened through Android Debug Bridge (ADB). In our current study, we selected 17 models, each running on the four chosen devices.
Figure 1: Overview of our research process. The rectangular boxes represent the procedures, while the ovals indicate the outcomes. We deployed 17 LLMs on four XR devices. First, we identified the key factors affecting performance. Based on these factors, we defined our evaluation criteria and conducted the experiments. Finally, we applied 3D Pareto analysis, which provided the Pareto front to highlight the best-performing model-device pairs. We also analyzed on-device LLMs versus cloud-based LLMs (highlighted in pink).
Due to compatibility constraints with Llama.cpp, we conducted testing exclusively on the CPU for Magic Leap 2, Meta Quest 3, and Vivo X100s Pro. As a smartphone, the Vivo X100s Pro theoretically supports GPU inference through frameworks like TensorFlow Lite or ONNX Runtime, but Llama.cpp does not currently offer GPU inference support for this device. Similarly, Magic Leap 2 and Meta Quest 3, being XR-specific devices, lack user-accessible GPU inference capabilities compatible with Llama.cpp. In contrast, Apple Vision Pro supports GPU inference via Metal, which we used in our experiments. These device and framework limitations necessitated restricting inference to CPUs for the first three devices.
3.2 Initial Investigation and Factors Identification
After deploying the LLMs, we conducted an initial investigation through anecdotal informal tests, literature reviews, and by studying LLM documentations. The aim of this preliminary investigation was to identify the key challenges and factors affecting performance—prerequisites for a fair evaluation—and to define our evaluation approach (Section 3.3).
We observed that XR devices experience performance degradation during prolonged tasks due to factors such as overheating, battery depletion, and background processes. Moreover, they often do not perform at a stable level. For instance, with the same model-device pair and parameters, we obtained different results across different runs of the same test. We refer to the metric used for measuring this variability as Performance Consistency. Another important factor is execution time: the time taken to execute a particular task. In our context, we used a relative term called Processing Speed, measured in tokens per second. We observed variations in processing speed depending on varying string length, batch size, and thread count. In addition, Memory Consumption and Battery Consumption are also important factors to consider. Similarly, we observed different processing speeds for different tasks. For example, prompt processing and token generation exhibited varying performance, even with the same model-device pairs and parameters.
In addition to the parameters mentioned above, during the initial investigation, we also noticed that ensuring consistent testing conditions across devices is critical for fair comparisons but challenging to maintain. For example, variance in a device’s Performance Consistency could create biases into the results. LLM performance may vary due to different architectures and device-specific optimizations. Even with identical parameters, the same device and LLM can produce different results due to unknown factors. The device’s environment, such as exposure to heat or obstructions to the cooling fan, can also affect performance. The concise presentation of the findings, and ensuring their reproducibility for re-evaluation and validation, is an additional challenge. Considering these challenges, we formulated a set of guidelines as our evaluation approach in Section 3.3.
3.3 Evaluation Approach
Considering the key challenges and our initial investigation (see Section 3.2), we formulate a standardized evaluation approach for our assessments. First, we designed an experimental setup to minimize potential biases and ensure the validity of the results (see Section 4). This setup establishes identical parameters across all devices and defines the test types to be executed. We developed several guidelines to ensure a comprehensive and standardized evaluation. To account for potential performance variations among LLMs, we selected 17 models from five different series (see Table I) and evaluated them across four devices. Additionally, we conducted Model Quality analysis to provide technical insights into the selected models. These tests are device-independent, and were conducted to understand the details and properties of each model. Next, each model-device pair was subjected to various tests (see Section 3.4).
To mitigate performance fluctuations, each experiment was repeated multiple times, and the average results were computed. Finally, we unified the five evaluation metrics, i.e., $[P_{1},P_{5}]$, using Pareto Efficiency theory to identify Pareto-optimal model-device pairs. Our evaluation framework and experimental design aim to facilitate reproducibility for other researchers. Additionally, we will provide open-source test scripts to enable replication of our study and further evaluations.
3.4 Models and Evaluation Metrics
We use 17 LLMs from five different series (see Table I), obtained from Hugging Face’s model repository [47] in our evaluations. For consistency in naming conventions, we assigned unique IDs ($m_{1}$ to $m_{17}$) to these models.
Table I: Models used, with their IDs, names, versions (v), quantization level (Q), layers (L), parameters (P), and sizes, sourced from Hugging Face [47].
Series ID Model Name v Q L P Size
Qwen $m_{1}$ qwen2-0_5b-instruct-fp16 v2.0 FP16 24 0.5B 0.942 GB
Vikhr-Gemma $m_{2}$ Vikhr-Gemma-2B-instruct-Q3_K_M v1.3 Q3 26 2.61B 1.36 GB
$m_{3}$ Vikhr-Gemma-2B-instruct-Q4_0 v1.3 Q4 24 2.61B 1.51 GB
$m_{4}$ Vikhr-Gemma-2B-instruct-Q5_0 v1.3 Q5 24 2.61B 1.75 GB
$m_{5}$ Vikhr-Gemma-2B-instruct-Q6_K v1.3 Q6 24 2.61B 2.00 GB
Phi-3.1 $m_{6}$ Phi-3.1-mini-4k-instruct-Q2_K v3.1 Q2 20 3.82B 1.32 GB
$m_{7}$ Phi-3.1-mini-4k-instruct-Q3_K_L v3.1 Q3 20 3.82B 1.94 GB
$m_{8}$ Phi-3.1-mini-4k-instruct-Q4_K_L v3.1 Q4 20 3.82B 2.30 GB
$m_{9}$ Phi-3.1-mini-4k-instruct-Q5_K_L v3.1 Q5 20 3.82B 2.68 GB
$m_{10}$ Phi-3.1-mini-4k-instruct-Q6_K v3.1 Q6 20 3.82B 2.92 GB
$m_{11}$ Phi-3.1-mini-4k-instruct-Q8_0 v3.1 Q8 20 3.82B 3.78 GB
LLaMA-2 $m_{12}$ llama-2-7b-chat.Q2_K v2.0 Q2 28 6.74B 2.63 GB
$m_{13}$ llama-2-7b-chat.Q3_K_S v2.0 Q3 28 6.74B 2.75 GB
Mistral-7B $m_{14}$ Mistral-7B-Instruct-v0.3.IQ1_M v0.3 IQ1 28 7.25B 1.64 GB
$m_{15}$ Mistral-7B-Instruct-v0.3.IQ2_XS v0.3 IQ2 28 7.25B 2.05 GB
$m_{16}$ Mistral-7B-Instruct-v0.3.IQ3_XS v0.3 IQ3 28 7.25B 2.81 GB
$m_{17}$ Mistral-7B-Instruct-v0.3.IQ4_XS v0.3 IQ4 28 7.25B 3.64 GB
Based on the Llama documentation [46] and our initial investigation, this evaluation comprises multiple tests to assess various performance parameters. For testing, we use Synthetic Prompts automatically generated by llama.cpp, ensuring uniform testing conditions without reliance on an external dataset. The string length is controlled via a parameter set to 64, 128, 256, 512, and 1024 tokens. The evaluation includes the following tests.
  • Performance Consistency: Measures stability over time.
  • Processing Speed: Analyzes performance across different string and prompt lengths, further divided into:
    • PP (Prompt Processing): Evaluates encoding efficiency in handling input prompts of varying lengths.
    • TG (Token Generation): Evaluates the device’s speed in generating output tokens of varying lengths.
  • Parallelization and Concurrency: This includes:
    • Batch Test (BT): Evaluates the device’s ability to process multiple input samples simultaneously by handling different batch sizes during token generation.
    • Thread Test (TT): Measures performance scalability by varying thread counts during LLM execution.
Here, Batch Size refers to the number of input samples processed in parallel. While larger batch sizes improve computational efficiency, they also demand more memory. Each of these metrics requires task-specific analysis for both models and devices. Furthermore, we conducted additional evaluations for Memory Usage and Battery Consumption. Finally, we performed a Pareto Optimality analysis to collectively examine the results across different models and devices.
4 Experimental Setup
This section describes the experimental setup and methodology for evaluating processing speed and error rates. The results are presented in Section 5. Before diving into the detailed setup, we first discuss the two evaluation metrics: processing speed and error count.
Processing Speed: In this study, we measure the processing speed for each model-device pair $(m_{i},d_{j})$ by recording the time $t_{ijk}$ taken by model $m_{i}$ running on device $d_{j}$ to process a string of length $L$. The instantaneous processing speed, $S_{inst.}$, expressed in tokens per second, is calculated as: $S_{inst.}=\frac{L}{t_{ij}}$. However, instead of relying on a single run, we perform 5 runs for each model-device pair to ensure accuracy and measure stability. If the coefficient of variance (CV) across the 5 runs is less than $33\%$, we calculate the 5-Run-Mean Speed (referred to as consistent speed, $S_{c}$) as:
$S_{c}=\frac{1}{5}\sum_{k=1}^{5}S_{inst.}(k),$ (1)
where $S_{\text{inst.}}(k)$ represents the instantaneous speed during the $k^{\text{th}}$ run of each experiment. This approach minimizes the impact of factors such as initialization delays, hardware fluctuations, or other environmental inconsistencies, providing a more reliable measure of processing speed.
Error Count: If the CV exceeds $33\%$, the run is considered unstable, and the corresponding data point is excluded from further analysis. The count of such unstable runs is recorded as an error during the experiment.
We conducted the following experiments.
4.1 Models Quality Analysis
Due to the storage and computational limitations of edge devices, deploying LLM locally often requires a trade-off between model size and performance. To put our performance results in perspective, we first conducted an empirical analysis of the models we have used in this study. This section offers insights into the models evaluated in this work, presenting a comparative analysis. While this section is not directly related to device performance, we believe providing detailed information about the models is essential for understanding their capabilities and limitations. We evaluate each model using six benchmarks, assessing various reasoning and knowledge capabilities: Hellaswag [48] (commonsense reasoning), MMLU [49] (multi-subject knowledge), ARC [50] (complex science questions), TruthfulQA [51] (truthfulness in responses), Winogrande [52] (pronoun resolution), and WikiText-2 [53] (language modeling and text generation).
For the first five benchmarks, which consist of multiple-choice questions, we use accuracy as the evaluation metric. However, for WikiText-2, the task is to predict the probability of each word in a given text, reflecting the model’s natural language understanding. Therefore, we use perplexity as the evaluation metric, defined as:
$\text{Perplexity}(P)=e^{-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i})},$ (2)
where $P(w_{i})$ represents the probability assigned by the model to the $i$-th word, and $N$ is the total number of words in the sequence.
4.2 Evaluating Performance Consistency
Some devices may experience performance degradation over time. In these experiments, we evaluate how processing speed fluctuates across multiple runs. For each model $m_{i}$ and device $d_{j}$, we perform $N$ runs and compute the mean $(\mu_{ij})$ of the speed $(S_{c})$, standard deviation $\sigma_{ij}$, and coefficient of variation ($CV_{ij}$). We also record the error count, as well as the maximum and minimum speeds observed during each experiment. In our experiments, we set $N=20$, i.e., each model-device pair is tested 20 times to evaluate performance variance. We noted the 5-Run-Mean speed $S_{c}$ measured during each run, calculated as in Eq. 1. In other words, each $S_{c}$ itself is the mean of 5 runs, ensuring that the measured value is not too affected by errors or hidden factors.
To let the devices cool down and conduct each run in comparable circumstances, we leave each device idle for two minutes between each run of the same model. We find this to be adequate, but just to be safer, between different models, we leave each device idle for ten minutes. For each test, we conducted both PP and TG evaluations with the string length set to 64, i.e., $\text{PP}=64$ and $\text{TG}=64$. For each model-device pair, we calculated the standard deviation, mean, and CV. The results are presented in Section 5.2.
4.3 Processing Speed and String Length
String length, comprising both Prompt Length and Size of the Token Set, significantly affects performance. We evaluate the performance of LLMs on XR devices with varying string lengths to assess their processing speed. For each experiment, the processing speed is calculated as in Eq. 1.
Figure 2: Model Quality Analysis: The models were evaluated based on six different benchmarks. We chose accuracy ($\uparrow$) for the five benchmarks on the left and perplexity ($\downarrow$) for the WikiText-2 benchmark. Each group in the bar chart represents models $m_{1}$ to $m_{17}$ from left to right.
We conducted experiments for five different string lengths (64, 128, 256, 512, and 1024) under the following PP and TG setups.
  • PP: The PP values were set to 64, 128, 256, 512, and 1024, while TG was fixed at 0, and other parameters at their default values.
  • TG: The TG values were set to 64, 128, 256, 512, and 1024, while PP was set 0, and other parameters were at default values.
Figure 3: Consistency results of the four devices over time: PP (top) and TG (bottom) speeds in tokens per second across 20 sorted runs (X-axis: run number, Y-axis: speed in $t/s$).
Figure 4: Consistency results of the AVP on GPU. Left: PP results. Right: TG results. Each model was tested 20 times, and the results are plotted in sorted order.
4.4 Parallelization with Thread Count and Batch Size
The parameters Batch Size and Thread Count, which control parallelization, also significantly affect performance. We evaluate the performance of LLMs on XR devices with varying batch sizes and thread counts to assess their processing speed, calculated using Eq. 1. The two tests conducted are described below:
  • BT: We set PP to 64, TG to 0, and varied the batch size (128, 256, 512, 1024), keeping other parameters at default values.
  • TT: We set PP to 64 and TG to 0, and varied the thread counts (1, 2, 4, 8, 16, 32), keeping other parameters at default values.
4.5 Evaluation of Memory Consumption
Memory consumption $M_{ij}$ is measured in terms of the Resident Set Size (RSS) which represents the actual physical memory usage of the relevant processes. For each model-device pair ($m_{i},d_{j}$), we measure memory usage over three runs, and report the average. The memory usage for PP and TG parameters was evaluated using the same approach as the performance consistency assessment. In each test, a fixed prompt was provided, and the model was instructed to generate outputs. The experiments were conducted with batch sizes of 128, 256, 512, and 1024, and their average memory consumption was recorded. The results are shown in Section 5.5.
4.6 Battery Consumption
This section presents battery consumption rates during extended PP and TG tests. From our initial investigation, we observed that only AVP exhibited a slight impact of model size on battery consumption, while the other three devices did not show significant variations with changing model sizes. Given this observation, instead of testing all 17 models, we conducted battery tests on selected models. Specifically, we chose $m_{1}$ as the sole model from the Qwen Series. For the remaining series, we selected the smallest and largest models: $m_{2}$ and $m_{5}$ from the Vikhr-Gemma Series, $m_{6}$ and $m_{11}$ from the Phi-3.1 Series, $m_{12}$ and $m_{13}$ from the LLaMA-2 Series, and $m_{14}$ and $m_{17}$ from the Mistral-7B Series. For each experiment, we recorded the battery level at the start and end of the experiment, allowing us to calculate battery consumption over a fixed duration of $600$ seconds (10 minutes). Each experiment was repeated three times, and we reported the mean values. To ensure consistent conditions, we took a $600$-second break between experiments, allowing the devices to cool down. The results are presented in Section 5.6.
4.7 Pareto Optimality
Although we evaluate our five performance metrics across various devices and models, the metrics remain fragmented and cannot be directly compared with other metrics. To address this, we turn to Pareto efficiency theory [54], which provides a measure of efficiency in multi-objective contexts and has been widely applied in various system design approaches [55, 56]. A choice is considered Pareto optimal if no other choice exists that can improve any of its objective criteria without deteriorating at least one other criterion. In our case, $x_{1}$ is considered dominated by $x_{2}$ through objects $f$ if :
$f_{i}(x_{1})\leq f_{i}(x_{2})\quad\forall i\in\{1,\dots,m\}\\ \text{and}\quad\exists j\in\{1,\dots,m\}\mid f_{j}(x_{1})<f_{j}(x_{2})$ (3)
where $i$, $j$ represent different objective indices.
A choice $x^{*}$ is considered Pareto optimal if no other feasible option dominates it. The set of all non-dominated designs forms the Pareto front, which represents the optimal trade-offs between all objectives.
To identify the optimal choices across devices and models, we define three objectives: quality, performance, and stability. To calculate the final score for each objective, we propose Eq. 4. For a given device-model pair and a specific objective, we first apply min-max normalization to each metric across all device-model pairs to eliminate the impact of different scales. Then, we compute a weighted sum of all metrics for that pair to obtain a single score:
$f_{o}(x)=\sum_{i=1}^{n}w_{i}\cdot\frac{m_{i}(x)-\min(m_{i})}{\max(m_{i})-\min(m_{i})},$ (4)
where $o$ is the objective index, $n$ is the number of metrics used in the objective category, $w_{i}$ defines the weight for each metric, and $\min$ and $\max$ calculate the minimum and maximum values across the entire set of results for metric $m_{i}$. In practice, some metrics, such as perplexity, are better when lower, which is the opposite of the direction in Eq. 3. Therefore, we take the reciprocal of each perplexity value.
Figure 5: Processing speed of the four devices in the PP test (top) and TG test (bottom) with varying string lengths: 64, 128, 256, 512, and 1024. The x-axis represents the model, while the y-axis represents the processing speed in t/s. For $m_{1}$, a different scale is used due to its higher values.
5 Performance Evaluation Results
This section presents the results of four types of experiments described in Section 4.
5.1 Model Quality Analysis
As shown in Table I, we select five different model architectures, and for each architecture, we choose various quantization settings. While lower-bit quantization reduces model size, speeds up inference, and lowers power consumption, it typically comes at the cost of reduced model performance. The results of the model quality are illustrated in Fig. 2. As shown in Fig. 2, a consistent trend is observed across all benchmarks: applying lower-bit quantization settings results in reduced model performance. Models with more parameters exhibit better language understanding capabilities, leading to higher performance under the same quantization settings. Additionally, different benchmarks exhibit varying levels of sensitivity to the quantization settings. For example, Hellaswag, ARC and WikiText-2 are more sensitive compared to the other three benchmarks.
5.2 Performance Consistency Results
Figs. 3 and 4 illustrate the consistency of performance across all model-device pairs. Across all 20 runs, the AVP demonstrates the most consistent results compared to other devices, with the corresponding speed also being the fastest. Notably, the GPU results for the AVP exhibit greater stability and smaller variance. Among the remaining devices, the Magic Leap 2 also shows a reasonable degree of stability; however, for the first three models, its variance is surprisingly high, reaching 27%, 32%, and 26% for $m_{2}$, $m_{3}$, and $m_{3}$ under the PP setup, and 20%, 23%, and 27% under the TG setup, respectively. In contrast, the Meta Quest 3 and Vivo X100 Pro exhibit relatively poor performance both in terms of speed and variance. Between these two devices, the Meta Quest 3 achieves higher speeds than the Vivo X100 Pro, but there is no clear difference in variance. For some models, one device shows lower variance, while for others, the variance is higher.
5.3 Results of Processing Speed and String Length
Fig. 5 presents the processing speed results of PP and TG with varying string lengths (prompt length and token sets). Both PP and TG were tested with values of 64, 128, 256, 512, and 1024. The processing speed varies significantly across string lengths, devices, and models.
String length and PP vs TG Speed: PP speed is consistently higher than TG speed, with a speedup of approximately two to three times across the four (CPU-based) devices, while on AVP (GPU), the PP speed is significantly higher, ranging from 10 to 19 times the TG speed. Similarly, increasing the prompt size slightly reduces the prompt processing speed in PP, whereas in TG, the speed reduction is more significant with a growing token set. This suggests that TG has a higher computational cost due to its sequential execution of tokens, whereas PP is somewhat parallelized, requiring only the loading of prompts. Though there is minor variation in the 64, 128, and 256 string lengths, the last two, 512 and 1024, are consistently the slowest. When examining the variation across string lengths, MQ3 exhibits the largest variance, whereas ML2 and AVP have the smallest variance.
Figure 6: Batch Test results: Batch size (X-axis) vs. processing speed (Y-axis).
Figure 7: Thread Test (TT) results: The X-axis represents the thread count, while the Y-axis shows the speed in tokens per second. Threads 4 and 8 deliver the fastest results. Note: AVP (CPU) fails for thread counts of 16 and 32.
Figure 8: Error counts for Meta Quest 3 (left) and AVP (right) in PP (top) and TG (bottom). Magic Leap 2, with zero errors, and Vivo X100 Pro, with a total of four errors, one in $m_{1}$ (PP-64), two in $m_{10}$ (PP-64, PP-128), and one in $m_{11}$ (TG-1024) are excluded.
Device-Based Analysis: If we examine the devices, the overall performance trends reveal that AVP consistently shows the highest speed in both PP and TG tests. ML2 ranks second in both PP and TG performance, while Vivo X100s Pro outperforms MQ3 in TG (particularly in longer strings 512, 1024) but ranks fourth in PP. This ranking highlights the varying computational capabilities of XR devices, with AVP (both CPU and GPU) standing out as the fasted one across all conditions. ML2 shows stable performance across all variations of string lengths, with lower $CV\%$ values, indicating that it does not vary significantly with changing string lengths. In contrast, MQ3 shows a higher $CV\%$, indicating greater variance with increasing string length.
Error Counts: Fig. 8 shows the errors counts during the PP and TG tests. Errors mostly occurred in AVP (CPU) and MQ3, particularly at longer prompt lengths of 512 and 1024 tokens. There are a few reasons for this. First, longer strings process more slowly, leading to abrupt changes and occasional failures. Second, memory constraints or inefficiencies in sustained generation contribute to these errors. With AVP, the general user experience is also suboptimal, as the device must be actively mounted on the head, which can cause errors if not handled carefully. Overall, MQ3 records the highest count of errors, with frequent retries affecting its reliability, whereas ML2 remains highly stable, leading to minimal inconsistencies (zero errors). As expected, AVP (GPU) also reported zero errors.
Model-Based Analysis: In the model-wise analysis, smaller models such as Qwen2-0.5B and Vikhr-Gemma-2B achieve the highest speeds across devices, particularly in PP. In contrast, larger models like the LLaMA-2-7B series and Mistral-7B series exhibit significantly lower processing speeds, with higher variability and instability, especially in TG. Model size also impacts processing speed, as smaller models are faster. Here, $m_{1}$ is the fastest (though omitted from later analysis), while $m_{2}$, $m_{3}$, and $m_{5}$ are the fastest among the remaining models, whereas $m_{14}$ and $m_{16}$ are the slowest.
5.4 Results of Parallelization with BT and TT
This section presents the results of BT and TT, both of which are used to achieve concurrency and parallelization. Regardless of the parallelization method, the three models $m_{2}$, $m_{3}$, and $m_{5}$ consistently rank among the fastest across all devices, while $m_{14}$ and $m_{16}$ are the slowest.
Results of BT: Fig. 6 presents the results of the batch test with varying batch sizes of 128, 256, 512, and 1024. Generally, increasing the batch size leads to a decrease in performance across all four devices (except for AVP (GPU)) due to increased computational overhead. AVP (GPU), with its strong computational resources, does not show a significant performance drop with varying batch sizes, although there are some fluctuations at a batch size of 256. This indicates that GPU-based processing on AVP (GPU) is highly optimized for parallel execution and can effectively manage larger batches of input without affecting processing speed.
More importantly, AVP (CPU) experiences a noticeable performance drop as the batch size increases. However, the remaining devices do not follow this trend. For instance, on MQ3, batch size 128 yields the best performance, whereas for the other batch sizes, there is no significant change. Similarly, for ML2 and Vivo, variations in batch size do not appear to have a substantial impact on performance.
Results of TT: Fig. 7 presents the TT results for thread counts of 1, 2, 4, 8, 16, and 32. Across all four devices, processing speed sees the most significant increase when increasing the thread count from 1 to 4. The speed gain continues at 8 threads (or shows minor degradation), but beyond 8 threads, performance begins to decline slightly, with further degradation beyond 16 threads. AVP (CPU) fails at thread counts of 16 and 32, indicating its limitations in achieving this level of parallelism.
AVP (GPU) follows a different trend, maintaining consistently high processing speed across all thread counts. Unlike CPU-based devices, its performance remains stable even at 32 threads, demonstrating its superior ability to handle concurrent tasks. These results highlight that while the four CPU-based devices benefit from moderate threading, AVP (GPU) is significantly more efficient at scaling concurrency without experiencing notable performance degradation. Since our study primarily focuses on CPU-based implementation, we conclude that using a moderate thread count of 4, 6, or 8 yields optimal results.
Figure 9: Memory consumption for each model-device pair. The results represent the mean values across batch sizes 128, 256, 512, and 1024.
5.5 Memory Consumption Results
Fig. 9 presents the memory consumption results for each model-device pair. The values represent the mean memory consumption from five experiments with varying batch sizes of 128, 256, 512, and 1024. As expected, memory consumption varies across models, with most showing consistent usage across different devices. Generally, within each model series, memory consumption increases with model size. However, minor exceptions exist, such as $m_{7}$ on AVP (GPU), which consumes more memory than other models in the Phi-3.1 series $[m_{6}-m_{11}]$. This size-based trend does not necessarily hold across series, even for models of similar sizes. For example, m2 (1.36 GB) consumes less memory than $m_{6}$ (1.32 GB) despite having a larger size, as they belong to different series. Similarly, $m_{7}$ (1.94 GB) requires more memory than $m_{5}$ (2.0 GB) and $m_{15}$ (2.05 GB), as all three belong to different series. Another example is $m_{12}$ (2.63 GB) consuming more memory than $m_{16}$ (2.81 GB), despite being smaller in size.
From a device-specific perspective, AVP demonstrates significantly lower memory consumption than the other devices, averaging $0.5782$ GB on the CPU and $0.5488$ GB on the GPU across all $17$ models. This efficiency is attributed to Apple Silicon’s unified memory architecture and the optimized memory management of the AVP, which can load only the active portions of a model during execution; for further details, see Feng et al. [57].
Among the remaining devices, Meta Quest 3 consumes the least memory at $2.3520$ GB, followed by Vivo X100 Pro at $2.3540$ GB, while Magic Leap 2 exhibits the highest at $2.3748$ GB. These findings suggest that AVP (both CPU and GPU) is approximately four times more memory-efficient than Magic Leap 2, Meta Quest 3, and Vivo X100 Pro. Its superior memory management and hardware optimizations make it the most efficient device in terms of memory consumption.
5.6 Battery Consumption Results
Fig. 10 presents the average battery consumption results over a 10-minute experiment, with each test repeated three times. We observe that for the first two series, the Qwen Series and Vikhr-Gemma Series, battery consumption remains relatively low and with lower variation across all four devices. However, for larger models in the LLaMA-2 Series and Mistral-7B Series, battery consumption increases significantly. AVP shows a strong correlation between model size and battery consumption, both for GPU and CPU usage. In contrast, the remaining three devices do not exhibit a significant variation with respect to model size. Vivo demonstrates the best battery life, with an average loss of only $2.5\%$ over 10 minutes. ML2 follows with a $8.5\%$ average battery loss, while MQ3 records a slightly higher battery loss at $9.7\%$. AVP shows the worst in this regard, with GPU usage resulting in a $10.1\%$ battery loss and with CPU usage leading to a $12.6\%$ battery loss in 10 minuets.
Figure 10: Battery consumption over 10-minute experiments, averaged across three runs.
Figure 11: Pareto fronts for various device-model pairs, with different colors representing distinct device types. Performance is evaluated based on processing speed, memory consumption, and battery consumption. Stability is assessed using the CV and error count values of PP and TG. Quality is measured through evaluation results on six benchmark datasets.
5.7 Pareto Front Results
Since we use both CPU and GPU for inference in the Vision Pro experiments, while the other three devices use CPU only for inference, we exclude AVP GPU from the Pareto calculation to ensure a fair comparison. With 4 devices and 16 models (m2-m17), we have a total of 64 device-model pairs. Using Eq. 3 and Eq. 4, we calculated the quality, stability and performance scores for each pair, then identified the Pareto fronts and visualized them in Fig. 11. The quality of a model can be evaluated using various benchmarks, such as accuracy and perplexity, as detailed in Section 4.1. For the performance objective, we evaluate performance using PP, TG, memory consumption, and battery consumption, with respective weights of 0.35, 0.35, 0.2, and 0.1. For the stability objective, we assess it based on the coefficient of variation (CV) and error count of PP and TG, assigning 0.7 to CV and 0.3 to error count.
Fig. 11 illustrates the Pareto front points, highlighting the optimal choices across different model sizes and devices. For example, the AVP (CPU) device achieves the highest speed among the models but exhibits relatively low stability, whereas the ML 2 device offers the greatest stability. Meanwhile, both Vivo and Meta Q3 deliver a balanced trade-off between speed and stability. At the model-pair level, $m_{1}7$ and $m_{1}1$ provide high quality on AVP (CPU) and Vivo, respectively. At the same time, $m_{3}$ is selected as a Pareto-optimal choice due to its high performance on AVP (CPU) and robust stability on Meta Q3. A similar trend is observed with $m_{8}$, which attains strong performance and quality on AVP (CPU) as well as commendable quality and stability on ML 2.
6 Analysis of On-Device LLMs
In this section, we evaluate the applicability of on-device LLMs in two interactive applications and report performance in terms of accuracy. Additionally, to assess efficiency, we compare on-device LLMs with alternative deployment approaches, namely client-server and cloud-based setups.
6.1 Evaluation on Interactive Applications
We conducted an analysis of our on-device LLM setup (AIvaluateXR) using datasets from two existing LLM-based interactive systems: VOICE [12] and GeoVis [58]. VOICE employs LLMs for interactive exploratory visualization and provides natural language interaction capabilities that allow users to navigate and manipulate 3D biological models in real time. GeoVis is another LLM-based method applied to geospatial data, enabling tasks such as selecting the viewing direction or retrieving location information through conversational queries. The details of the metrics, prompts and queries used are provided in the supplementary material. For this evaluation, we used the following metrics:
  • Formatting (GeoVis): Measures how well the model’s output conforms to the expected [latitude, longitude] format. A score of 100% is assigned if the formatting is perfect, 75% if it is off by one character, 50% if it is otherwise incorrect but still easy to parse, and 0% if it cannot be parsed.
  • Accuracy (GeoVis): Evaluates the spatial accuracy of the predicted coordinates relative to the target location. The score is 100% if the coordinates are near the center of the target area, 20% if they are off target but the area of interest remains partially visible at a reasonable zoom level, and 0% if it is not visible at all.
  • Score: For the VOICE dataset, each query is graded 100% if the output is correct and 0% otherwise. For the GeoVis dataset, if the Formatting score is 0, the overall Score is also 0. Otherwise, it is computed as a weighted combination of Formatting (20%) and Accuracy (80%) to reflect both structural and spatial correctness.
Since our current study focuses specifically on LLMs, we did not take evaluate the LLMs on prompts requiring a vision model or relying on fine-tuned models for modeling tasks. Instead, we used the VOICE and GeoVis datasets and evaluated only the queries that LLMs could answer on their own, which is sufficient for evaluating their accuracy. Table II presents the accuracy results. Since these tasks involve complex interactions, we observe that some models completely fail in terms of accuracy.
Table II: Performance summary of models $m_{8}$ to $m_{17}$ on the GeoVis [58] and VOICE [12] datasets. Each value $(\mu,\sigma)$ represents the mean ($\mu$) and standard deviation ($\sigma$), separated by a comma. The Score metric is computed as a weighted sum of Formatting (20%) and Accuracy (80%).
Model ID GeoVis data VOICE data
Formatting ($\mu,\sigma$) Accuracy ($\mu,\sigma$) Score ($\mu,\sigma$) Accuracy ($\mu,\sigma$)
$m_{8}$ 40.00 , 20.34 55.00 , 44.70 52.00 , 38.44 0.00 , 0.00
$m_{9}$ 40.00 , 20.34 57.33 , 46.23 51.20 , 39.85 0.00 , 0.00
$m_{10}$ 36.67 , 22.49 71.00 , 41.63 56.13 , 40.13 3.70 , 19.25
$m_{11}$ 45.00 , 15.26 64.00 , 45.00 60.20 , 37.57 0.00 , 0.00
$m_{12}$ 0.00 , 0.00 0.00 , 0.00 0.00 , 0.00 3.70 , 19.25
$m_{13}$ 1.67 , 9.13 0.00 , 0.00 0.33 , 1.83 0.00 , 0.00
$m_{14}$ 38.33 , 21.51 6.00 , 15.89 11.93 , 14.03 7.41 , 26.69
$m_{15}$ 52.50 , 7.63 39.33 , 42.74 41.97 , 34.92 48.15 , 50.92
$m_{16}$ 13.33 , 26.04 17.33 , 36.29 16.53 , 33.70 40.74 , 50.07
$m_{17}$ 54.17 , 13.27 87.33 , 29.47 80.70 , 23.63 51.85 , 50.92
GPT-4o 100 , 0 99.66 , 1.82 99.73 , 1.43 96.30 , 19.25
GPT-4o-Mini 98.33 , 9.13 93.33 , 22.94 94.33 , 19.79 85.19 , 36.20
Table III: Time comparison (in seconds), including both model loading (warm-up) time and first response latency, for model $M_{15}$ running locally on XR devices. Each cell reports the mean and standard deviation $(\mu,\sigma)$.
Model ID Series AVP ML2 Vivo MQ3
$m_{15}$ Mistral-7B 2.22 , 1.05 61.21 , 38.04 86.69 , 53.84 82.96,53.36
Table IV: Comparison of cold and warm first response time (F.R.T) in seconds for seven LLMs accessed via AVP across three deployment modes. Each cell reports the mean and standard deviation $(\mu,\sigma)$. Here, $\times$ indicates that the model is not supported in that mode.
Type Approach GPT-4o GPT-4o Mini Qwen2 ($m_{1}$) Vikhr-Gemma ($m_{5}$) Phi-3.1 ($m_{11}$) LLaMA-2 ($m_{13}$) Mistral ($m_{17}$)
Cold F.R.T Client-Server $\times$ $\times$ 0.67, 0.66 1.94, 1.36 2.22, 1.05 2.65, 1.54 3.72, 0.71
On-Device $\times$ $\times$ 1.07, 0.06 2.56, 0.37 3.06, 1.17 3.30, 0.93 3.86, 0.68
Warm F.R.T Cloud-based 1.13, 0.29 1.03, 0.31 $\times$ $\times$ $\times$ $\times$ $\times$
Client-Server $\times$ $\times$ 0.25, 0.15 1.36, 0.69 1.37, 0.07 1.40, 0.23 2.91, 0.68
On-Device $\times$ $\times$ 0.03, 0.00 0.26, 0.28 0.12, 0.04 0.27, 0.28 0.07, 0.00
6.2 Latency Comparison Across Deployment Modes
Latency, or first response time (F.R.T), refers to the duration between issuing a prompt and receiving the first token of the response, whereas warm-up time denotes the initial time required to load the model into memory. Cold F.R.T refers to the latency observed on the first prompt after model initialization (i.e., warm-up $+$ response), while Warm F.R.T reflects the prompt-to-first-token latency for subsequent queries on an already loaded model. We compared latency across three deployment modes: cloud-based, client-server, and on-device. In the client-server setup, the LLMs were hosted on a MacBook Pro with an M4 Pro chip and 48 GB of RAM, and accessed via AVP. To begin, we evaluated model $m_{15}$ across all four XR devices (see Table III) and selected the AVP as the fastest device for subsequent comparisons. For a broader evaluation, we selected five representative models (one per category) and benchmarked them across all three deployment modes. Each experiment was repeated three times, and we report the mean latency along with standard deviations.
As shown in Table IV, these five models were evaluated alongside GPT-4o and GPT-4o-mini, all accessed via the AVP. While on-device models exhibit longer warm-up times (i.e., higher Cold F.R.T), they consistently achieve the lowest latency after loading (i.e., lower Warm F.R.T), making them well-suited for real-time XR interactions, since warm-up only happens once during a session. Importantly, the latency values in Table IV correspond to a simple prompt—“Give me 20 continuous words as text only”. For longer or more complex prompts, the delay can further increase, potentially rendering cloud-based models unsuitable for time-sensitive XR applications.
[…]
Furthermore, over time, XR devices will feature more powerful hardware, and perhaps dedicated AI accelerators, so the latencies of on-device LLMs will diminish. On the other hand, cloud-based LLMs may gradually process prompts faster for the same reasons, but there will always be some network latency that is unlikely to shrink substantially. The relatively small difference in latency that we observed between GPT-4o and GPT-4o-mini suggests that network latency may already dominate overall latency for these models, indicating that this overall latency is unlikely to shrink by much [59]. Therefore, over time, the first-response time advantage that on-device LLMs currently enjoy is likely to grow, while their warm-up time disadvantage is likely to shrink, so long as the LLMs are small enough to be run efficiently on XR devices.
7 Discussion
AIvaluateXR provides a systematic framework for evaluating LLMs running locally on XR devices, taking into account both model- and device-specific factors. Using AIvaluateXR, we have performed extensive testing of multiple open-source language models on four distinct XR devices. Our study identifies critical performance evaluation factors and introduces a suite of testing scripts to analyze these aspects systematically. These scripts enhance the reproducibility of our findings, making them relevant for future analyses in similar contexts. The processing speed, measured in tokens per second, was evaluated across multiple dimensions. This includes a consistency analysis to determine whether each model-device pair maintains stable performance over time or exhibits fluctuations. The impact of string length on both PP and TG was analyzed for each model-device combination. Additionally, the study examined the support for parallelization and concurrency control through BT and TT tests. Furthermore, memory usage and battery consumption were also assessed to provide a holistic evaluation of system performance.
7.1 Summary of Results
Figs. 4 and 3 present the consistency results in ascending order, where we see that the AVP GPU exhibits the highest stability. While ML2 shows a similar level of stability to the AVP CPU, its first three models ($m_{2}$, $m_{3}$, and $m_{4}$) exhibit a CV above $20\%$ for TG and above $26\%$ for PP. Considering the error count in Fig. 8, ML2 emerges as the most stable device, having recorded zero errors. In contrast, the AVP CPU encountered multiple errors. The remaining two devices, Vivo and MQ3, exhibit higher instability, with Vivo performing relatively better than MQ3. This also indicates that consistency is model-dependent, as seen with models $m_{2}$, $m_{3}$, and $m_{4}$, which exhibited reduced stability.
For $m_{1}$, being the smallest model, all devices performed efficiently. Among the remaining models ($m_{2}$ to $m_{16}$), the average inference speeds across various devices (according to PP and TG tests) are as follows. AVP (CPU) achieved the highest inference speed (19.91 t/s PP, 11.54 t/s TG), followed by ML2 (12.18 t/s PP, 6.73 t/s TG), MQ3 (10.89 t/s PP, 5.17 t/s TG), and Vivo (9.35 t/s PP, 6.12 t/s TG), reflecting performance variations across XR devices. These speeds meet basic conversational requirements and enable LLM inference for broader applications beyond dialogue-based tasks. The PP speed across all devices is two to three times faster than TG speed, while for AVP GPU PP speed is 10 to 19 time faster than TG speed. This is because PP only encodes the input, whereas TG must process sequential dependencies during decoding. In model-based analysis, smaller models such as $m_{1}$, $m_{2}$, $m_{3}$, and $m_{5}$ are the fastest, while $m_{14}$ and $m_{16}$ were the slowest. Regarding string length, larger PP and TG inputs (512, 1024) are slower than shorter ones (64, 128, 256). Moreover, longer strings (512, 1024) frequently result in errors. Among devices, MQ3 exhibited the highest number of errors, followed by the AVP CPU and then Vivo. In contrast, ML2 and the AVP GPU did not experience any significant errors or inconsistencies.
In batch tests (BT), performance generally decreases as batch size increases across all devices, except for the AVP GPU, which remains stable regardless of batch size. The AVP CPU, however, experiences a strong performance degradation as batch size increases, more pronounced than in other devices. For thread tests (TT), across all four devices, mid-range thread counts (4 to 8) yielded the best performance, while lower (1, 2) and higher (16, 32) thread counts resulted in lower performance. Notably, the AVP CPU fails at thread counts of 16 and 32, whereas the AVP GPU exhibits its best performance at these higher thread counts.
In terms of memory consumption, significant variations were observed across different models. Generally, larger models consumed more memory. Interestingly, each of the five model series displayed distinct memory consumption trends. Some smaller models from one series consumed more memory than larger models from another series. For instance, $m_{2}$ (with size 1.36 GB) consumes less memory than $m_{6}$ (with size 1.32 GB), highlighting the model-dependent nature of memory consumption (see Fig. 9).
Battery consumption did not show significant differences across devices, although the AVP exhibited slightly higher battery life loss, with an average battery depletion of up to $10.1\%$ (with GPU) and $12.6\%$ (with CPU) over 10 minutes. In contrast, the Vivo X100 Pro performed the best, with only a $2.5\%$ loss, while the Magic Leap 2 demonstrated good battery efficiency with an $8.5\%$ loss. The Meta Quest 3 consumed an average of $9.7\%$ over the same duration.
With 3D Pareto Optimality, we found AVP leading overall, ML2 demonstrating high consistency and stability, Vivo being slower but still consistent in performance, and MQ3 ranking lowest across all aspects. Among the models, $m_{3}$, and $m_{5}$ performed particularly well, with smaller models generally exhibiting better results. In contrast, models from the LLaMA-2 and Mistral-7B series were the slowest. We have analyzed the performance of an on-device LLM using 57 queries from an interactive application dataset. Our findings suggest that with targeted training and fine-tuning, the performance can be further improved.
7.2 The Impact of LLM Characteristics
The size of the model on disk, which correlates with the number of parameters, has a substantial impact on performance. In this study, we primarily selected models below 4 GB in size to ensure they could be deployed on all tested devices. We observed that larger models not only exhibited lower inference speeds and higher memory usage but also consumed more battery compared to smaller ones. However, in the context of interactive application accuracy, larger models such as the Mistral-7B and Phi-3.1 series achieved higher accuracy, while LLaMA-2 performed comparatively poorly. It is important to note that our testing for interactive applications was conducted without any fine-tuning. This suggests that selecting a smaller model—for example, from the Qwen series—and fine-tuning it on XR-specific data could achieve a better balance of accuracy and efficiency. Overall, these findings indicate the need to carefully balance model size and parameter count to reach optimal trade-offs between speed, memory consumption, battery life, and accuracy.
7.3 The Impact of Device Capabilities
We found that the AVP achieved the highest speed in both PP and TG tests. ML2 ranked second in overall performance, while the Vivo X100s Pro outperformed MQ3 in TG performance but ranked fourth in PP speed. These results can be attributed to the AVP’s unified memory architecture and hardware-accelerated memory management, which result in lower memory consumption and more stable performance. Furthermore, the combination of hardware design and operating system-level memory management plays a critical role in determining device efficiency. The remaining three devices demonstrated comparatively weaker performance, underscoring that selecting an XR device for AI workloads depends heavily on available memory and the underlying memory management capabilities. While understanding these factors beforehand is beneficial, applying AIvaluateXR provides a systematic way to quickly assess and compare device suitability for AI applications in XR environments.
7.4 Implications for XR Applications
Based on our evaluation of XR interactive applications, specifically VOICE and GeoVis (as shown in Table II), we observe that even without any domain-specific training, larger models achieved reasonably good accuracy. While smaller models demonstrated lower accuracy, their performance could likely be improved through grounding or fine-tuning on XR-specific datasets. The trade-offs between accuracy and processing time linked to model size are an important consideration for practical deployment but can be mitigated through targeted domain adaptation. These findings highlight the potential of AIvaluateXR as a practical tool for identifying and optimizing model-device configurations in this emerging area at the intersection of AI and XR.
7.5 The Future of On-Device LLMs
While cloud-based LLMs offer powerful capabilities, they may not always be preferred due to factors such as limited internet connectivity, subscription costs, privacy concerns, or regulatory restrictions. An alternative approach is the client-server model, where LLMs run on a local server and XR devices connect over a local network. However, this setup still introduces network latency (see Table IV) and does not fully address data security risks. By contrast, running LLMs directly on-device makes the system more autonomous and independent of external infrastructure. Looking ahead, we anticipate that advances in model compression, quantization techniques, and specialized hardware accelerators will further reduce resource demands and latency, making fully on-device intelligence increasingly practical. Given the rapid growth in XR and wearable devices, enabling on-device LLMs has the potential to unlock a wide range of real-world applications across domains, from education and training to healthcare, industrial maintenance, and beyond.
7.6 Limitations
Apart from AVP, our study focuses on CPU-based evaluation, as limited GPU memory on the three devices hindered full model loading, affecting performance. Finally, since the study is primarily based on Llama.cpp implementations, its generalizability is limited.
8 Conclusion
AIvaluateXR provides a comprehensive framework for deploying LLMs on XR devices and conducting in-depth evaluations across multiple performance factors. Using this framework, we benchmarked 17 LLMs across four XR devices to assess their performance in a systematic and reproducible manner. Given the GPU memory limitations of most XR devices, our study primarily focuses on CPU-based analysis. However, since the AVP has sufficient GPU resources, we also include GPU results for AVP, providing additional insights into the study. Generally, for processing speed, we can rank the four devices in the following order: AVP, ML2, Vivo, and MQ3, although MQ3 outperforms Vivo in TG.
In memory consumption, AVP was found the most memory-efficient. Among the remaining devices, differences were minimal, though ML2 occasionally consumed the most memory. As expected, larger models required more memory, but this trend varied across different model series. For BT tests, a lower batch size (128) yielded the best performance, while larger batch sizes reduced speed for CPU-based tests, with AVP GPU remaining largely unaffected. For TT tests, medium thread counts (4–8) were optimal across all devices. In terms of consistency, ML2 is the most stable, exhibiting the lowest variance in speed and zero errors. AVP GPU also maintains low variance. MQ3 performs the worst in this aspect, while Vivo is more stable than MQ3 but less consistent than the rest. The Pareto analysis highlighted several models on AVP due to their outstanding performance, while the Pareto fronts found on ML2 because of their consistently stable results. Additionally, a few models on Vivo and MQ3 emerged on the Pareto front because of their balanced trade-offs between performance and stability. Models $m_{12}$ to $m_{17}$ were found slower.
Our latency comparison across cloud-based, client-server, and on-device deployments—along with two interactive application tests—demonstrates that while on-device LLMs currently yield slightly lower accuracy, this can be improved through domain-specific fine-tuning. Notably, despite longer warm-up times (i.e., higher Cold F.R.T), on-device models offer excellent latency post-loading (i.e., lower Warm F.R.T), making them highly promising for real-time XR applications. Given the rise of XR and LLMs, we believe on-device AI has strong potential, in which AIvaluateXR can play a key role.
Acknowledgments
This research has been funded by KAUST Competitive Research Grants ORFS-CRG12-2024-6422. We thank Prof. Kiyoshi Kiyokawa from NAIST, Japan, for valuable discussions. We also thank Deng Luo and Da Li from our team at KAUST for their support and valuable input across various modules.